OpenStudy (lilai3):

Please help! WILL GIVE A MEDAL AND IF YOU DON'T GET ONE, I'LL BE SURE TO FAN YOU!!!! (: The bicycle path at the park is a loop that covers a distance of m miles. Jorge biked 2 loops each on Monday and Wednesday and 3 loops on Friday. On Sunday, Jorge biked 10 miles. Write an expression in simplest form that represents the total distance that Jorge biked this week.

OpenStudy (compassionate):

m = 2 + 3 + 10 m = 15

OpenStudy (lilai3):

it's supposed to be an expression, not an equation

OpenStudy (compassionate):

2 + 3 + 10

OpenStudy (helpsos):

10+ 2(2m)+ 3m since you don't know how many miles the loops are you have to write it represent it by writting m. What do you think?

OpenStudy (lilai3):

is it 2m + 3m + 10 m =15m?

OpenStudy (helpsos):

see becuase he did 2 loops on monday and wednesday it has to be 2m+2m which can also be written as 2(2m). And then he did 3 loops on friday which is 3m. Then in the end you just add 10 miles.

OpenStudy (lilai3):

so..i'm right?

OpenStudy (helpsos):

umm give me a sec :)

OpenStudy (lilai3):

kk

OpenStudy (helpsos):

\[10+2(2m)+3m= total miles\] \[10+ 4m +3m= total miles\] \[10+ 7m= total miles \] this is what i got :) in simplest form

OpenStudy (helpsos):

does it make sense?

OpenStudy (lilai3):

sort of yes thank you. will you also be willing to help me on one more problem? ur so kind.

OpenStudy (helpsos):

np and you dont have to give medals or anything

OpenStudy (helpsos):

Id be happy to help I hope I get the problem though

OpenStudy (lilai3):

Mala is y years old. Her sister is 4 years older than Mala. Write an expression in simplest form that represents the sum of the ages of the sisters.

OpenStudy (helpsos):

ok so: Mala's age= y Sister's age = 4+y therefore the equation is: Mala+sister= sum of age y+(4+y)= sum of age does this make sense or is there anything confusing?

OpenStudy (helpsos):

if you want an expression you could just write: y+(4+y)

OpenStudy (lilai3):

oh no i'm checking my work on these

OpenStudy (lilai3):

thank you soo very much. :)

OpenStudy (helpsos):

np btw what grade are you in? these problems seems similar to something i'm learning in class write now :)

OpenStudy (lilai3):

uhm middle of middle school.

OpenStudy (helpsos):

oh i knew it! im in grade 9

OpenStudy (helpsos):

ure in grade 8?

OpenStudy (lilai3):

holy fudge! no 7th

OpenStudy (helpsos):

oh lol i was close :D

OpenStudy (lilai3):

yes.

OpenStudy (lilai3):

i got to go. bye. thanks

OpenStudy (helpsos):

well good luck with the rest of ure work and byeee :)